How much energy is stored in a 50pF capacitor when it is charged to a potential difference of 200V?
1
Expert's answer
2022-02-28T11:51:50-0500
"W=\frac{CV^2}{2}\\
=\rm\frac{50*10^{-12}\: F*(200\: V)^2}{2}=1.0*10^{-6}\: J=1.0\: \mu J"